Find the exact value of the trig function cos (u+v) given that sinu= 5/13 and cosu= -3/5
Answer by stanbon(75887)

Find the exact value of the trig function cos (u+v) given that sinu= 5/13 and cosu= -3/5
-----
Note:: You need to have some information about "v".
Cheers,
Stan H.